La presentazione è in caricamento. Aspetta per favore

La presentazione è in caricamento. Aspetta per favore

LHCf Status Report Measurement of Photons and Neutral Pions in the Very Forward Region of LHC Oscar Adriani INFN Sezione di Firenze - Dipartimento di Fisica.

Presentazioni simili


Presentazione sul tema: "LHCf Status Report Measurement of Photons and Neutral Pions in the Very Forward Region of LHC Oscar Adriani INFN Sezione di Firenze - Dipartimento di Fisica."— Transcript della presentazione:

1 LHCf Status Report Measurement of Photons and Neutral Pions in the Very Forward Region of LHC Oscar Adriani INFN Sezione di Firenze - Dipartimento di Fisica dellUniversità degli Studi di Firenze Frascati, CSN1, 26 Giugno 2007

2 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report INTERACTION POINT Beam line Detector II TungstenScintillator Silicon strips Detector I TungstenScintillator Scintillating fibers 140 m Experimental Method: 2 independent detectors on both sides of IP The direct measurement of the production cross section as function of p T is essential to correctly estimate the energy of the primary cosmic rays (LHC: eV) Simulation of an atmospheric shower due to a eV proton.

3 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Detector # 2 4 pairs of silicon microstrip layers (6, 12, 30, 42 r.l.) for tracking purpose (X and Y) impact point We use LHC style electronics and readout 16 scintillator layers (3 mm thick) Trigger and energy profile measurements Absorber 22 tungsten layers 7mm thick 44 X 0 (1.6 I ) in total (W: X 0 = 3.5mm, R M = 9mm) 2 towers 24 cm long stacked on their edges and offset from one another Lower:2.5 cm x 2.5 cm Upper: 3.2 cm x 3.2 cm Energy < 400 rad < 400 rad

4 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Performances of the LHCf Detector SciFi Position Resolution Energy Resolution Measured at the SPS Beam Test in 2004 LHCf can measure (and provide to LHC) the center of neutral flux from the collisions If the center of the neutral flux hits LHCf << 1 mm resolution << 1 mm resolution

5 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report 0 mass resolution 0 mass resolution Arm #1 E/E=5% E/E=5% 200 m spatial resolution m/m = 5% m/m = 5% 2 identified in the 2 towers

6 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Monte Carlo ray energy spectrum (5% Energy resolution is taken into account) (5% Energy resolution is taken into account) 10 6 generated LHC interactions 10 6 generated LHC interactions 1 minute 29 cm -2 s -1 luminosity Discrimination between various models is feasible Quantitative discrimination with the help of a properly defined 2 discriminating variable based on the spectrum shape (see TDR for details)

7 Where are we with the detectors?

8 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Where we are with Arm1? Arm1 was fully assembled in Japan in July 2006 (scintillators + fibers + Tungsten) and fully tested at CERN in August 2006 beam test

9 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Where we are with Arm2? o Arm2 was partially assembled in Florence in July 2006 and brought to CERN for the Beam Test of August 2006 o Arm2 was fully assembled in Florence in April 2007

10 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Front Counter (Both for Arm1&2) read by a MAPMT R7600U-00-M4

11 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report The mechanics of the module Silicon sensor Hybrid circuit Pace Chips Kapton fanout SiliconX SiliconY Tungsten Light Guides + Scintillators 25 mm

12 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report

13 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report

14 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Readout electronics LHC style electronics (Many thanks to CMS!!!!) DOHCCUPLLDCU… PACE3 chips (CMS preshower) We have developed Hybrid circuit (Left+Right)Hybrid circuit (Left+Right) Mother Board (Logic)Mother Board (Logic) Adc Piggy Back boardAdc Piggy Back board DAQ: 100 Mbit/s optical link VME Readout 100MbitOpticalLink

15 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Now the real installation in LHC… Pre-Installation dates have been fixed in Fall 2006 –LSS1L (Arm1) o 8/01/2007 to 26/01/2007 FINISHED! –LSS1R (Arm2) o 23/04/2007 to 11/05/2007 FINISHED The dates for the final installation are still under discussion (after the beam test in August/September, may be beginning of 2008?) The idea is that installation takes place in 2 steps: Pre-installation Final installation In between the 2 installation the baking out of the beam pipe will be done (200 o C), so the detectors should be removed

16 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Before pre-installation a long preparatory job has been necessary: –Modifications of the TAN –Transport machine / Fork lift –Cables and fibers from USA15 to tunnel o Most of the control electronics and all the power supply system are located in the USA15 Atlas control room o > 200 m distance!!! o All the necessary cables have been pulled (~100 cables!!!) o All the connectors have been mounted o All the optical fibers have been installed –Electronic racks in USA15

17 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report TAN slot before and after CU bar removal

18 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Arm1 pre-installation From 8/01/2007 to 26/01/2007 No major problems came out Cables OK Transport and installation OK Laser calibration OK Power supply from USA15 OK Manipulator and movements OK Arm1 was dismounted at the end

19 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Transport and insertion in the TAN

20 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report LHCf Arm 1 – Installation completed within 15 minutes!

21 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Arm2 pre-installation From 23/04/2007 to 11/05/2007 No major problems came out Quartz fibre for calibration is damaged should be replaced Cables OK Transport and installation OK Laser calibration OK locally TBD from remote TBD from remote Power supply from USA15 OK Manipulator and movements OK No interference with BRAN Arm2 was dismounted at the end Explicit request from TS/LEA: Speed up the cabling connection time! Some work is in progress for this (Splitter box, Distribution board) A second pre-installation is foreseen in Autumn

22 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report

23 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report TANUSA15 High Current Cables Sense wires Power Supply High Voltage Splitter Box ~ 240 m Thermistors Control PC MB + ADC TSC+FEC VME Crate Monomode Optical Fibers DOH Multimode Optical Fibers Distribution Board Power Thermistors Signal + Control Arm2 Readout and control scheme for Arm2 Silicon

24 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report General LHCf DAQ scheme LHCf Rack Allocation in USA15

25 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report What do we have to do? Final commissioning of the electronics (modifications to the logic, complete the PCBs, distribution board etc.) Final commissioning of the electronics (modifications to the logic, complete the PCBs, distribution board etc.) Installation of the new splitter box in the tunnel Installation of the new splitter box in the tunnel Some work in USA15 and TAN region to speed up the interconnection time Some work in USA15 and TAN region to speed up the interconnection time Final DAQ software common with Japan people Final DAQ software common with Japan people Monitoring program Monitoring program Test beam with final apparatus (August 24, September 11) Test beam with final apparatus (August 24, September 11) Second pre-installation to test the modifications done after the first pre-installation Second pre-installation to test the modifications done after the first pre-installation Final installation: schedule is under discussion (December? January 2008?) Final installation: schedule is under discussion (December? January 2008?)

26 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Milestones dei preventivi /02/ /02/2007 –completamento dei moduli al silicio FATTO! 30/06/ /06/2007 –Preinstallazione e commissioning di LHCf nella TAN compatibilmente con le tempistiche della macchina FATTO! 31/12/ /12/2007 –Installazione finale e presa dati a bassa energia compatibilmente con la macchina –Installazione finale e presa dati a bassa energia compatibilmente con la macchina Installazione prevista per la fine dellanno (inizio 2008?) Installazione prevista per la fine dellanno (inizio 2008?) presa dati sicuramente NO! presa dati sicuramente NO!

27 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report 2007 Test Beam 3 settimane di Test Beam (24/08 – 11/09) 3 settimane di Test Beam (24/08 – 11/09) SPS H4 SPS H4 Scopo: Scopo: –Calibrazione assoluta in energia di Arm1 & Arm2 utilizzando più energie possibili per gli elettroni (50 GeV – 250 GeV) –MIsura di E/E per protoni –Misura congiunta LHCf-BRAN (ultimi giorni di test LHCf – Primi giorni di test BRAN) –MIsura diretta di produzione 0 con target leggero?

28 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status ReportMoney…. Per questo lascio la parola ai referee… Ma prima un breve commento: Principale problema riscontrato per il controllo dei costi: Mancanza di stime precise per le spese necessarie per linstallazione. I lavori vengono svolti da ditte esterne sotto il controllo della divisione TS I costi preventivati normalmente non vengono rispettati (difficoltà nellinstallazione, allarmi di evacuazione nel tunnel, extra costi ….) Continua negoziazione per laccettazione e la ripartizione delle spese tra i vari utenti (Atlas, LHCf, CERN) 2006: Problema del costo aggiuntivo dei cavi 2007: Problema del costo aggiuntivo delle fibre 2008: Speriamo nessun problema!!!! Necessità di un accordo ufficiale con Atlas e il Cern per le spese per gli anni futuri?

29 Back up Slides

30 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Detector #1 16 scintillator layers (3 mm thick) Trigger and energy profile measurements Absorber 22 tungsten layers 7mm thick 44 X 0 (1.6 I ) in total (W: X 0 = 3.5mm, R M = 9mm) 4 pairs of scintillating fiber layers for tracking purpose (6, 10, 30, 42 r.l.) Energy Impact point ( ) 2 towers ~24 cm long stacked vertically with 5 mm gap Lower:2 cm x 2 cm area Upper: 4 cm x 4 cm area < 300 rad

31 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Energy spectrum of π 0 expected from different models (Typical energy resolution of is 3 % at 1TeV)

32 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Clear CR signal separation (20ADC by 2249W ~ 5pC)

33 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Model dependence of neutron energy distribution Original n energy 30% energy resolution

34 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Installation of the detectors in the TAN absorbers at 140m from IP1 96 mm LHCf

35 The LHCf operation in LHC

36 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report Optimal LHCf run conditions Beam parameters used for commissioning are good for LHCf!!! Beam parameter Value # of bunches Bunch separation > 2 sec Crossing angle 0 rad 0 rad 140 rad downward Luminosity per bunch < 2 x cm -2 s -1 < 2 x cm -2 s -1 Luminosity ~10 30 cm -2 s -1 ~10 30 cm -2 s -1 Bunch intensity 4x10 10 ppb ( *=18 m) 1x10 10 ppb ( *= 1 m) No radiation problem for 10kGy by a year operation with this luminosity

37 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report From R. Bailey presentation at January 2007 TAN workshop Stage 1 Optimal conditions for LHCf running!

38 Oscar Adriani Frascati, 26 Giugno CSN1 LHCf Status Report LHCf proposed running scenario Phase-I Phase-I Run since the very beginning of LHC operations (Stage 1, 43 bunches) Run since the very beginning of LHC operations (Stage 1, 43 bunches) Remove the detector for radiation issues when the machine goes to the Stage II (luminosity reaches cm -2 s -1) and reinstall the 3 Cu bars Remove the detector for radiation issues when the machine goes to the Stage II (luminosity reaches cm -2 s -1) and reinstall the 3 Cu bars Phase-II Phase-II Re-install the detector at the next opportunity of low luminosity run after removal of Cu bars (Totem dedicated runs? Possible LHCf dedicated runs?) Re-install the detector at the next opportunity of low luminosity run after removal of Cu bars (Totem dedicated runs? Possible LHCf dedicated runs?) Phase-III Phase-III Future extension for p-A, A-A run with upgraded detectors? Future extension for p-A, A-A run with upgraded detectors? We are continously in touch with Machine and ATLAS peoples to agree on detailed running scenario


Scaricare ppt "LHCf Status Report Measurement of Photons and Neutral Pions in the Very Forward Region of LHC Oscar Adriani INFN Sezione di Firenze - Dipartimento di Fisica."

Presentazioni simili


Annunci Google